OpenStudy (anonymous):

Choose the missing step in the given solution to the inequality -6x - 10 > 14 + 2x. -6x - 10 > 14 + 2x -8x > 24 x < -3

OpenStudy (anonymous):

do you have answer choices ?

OpenStudy (anonymous):

yes , -4x - 10 > 14 -8x - 10 > 14 -4x - 10 < 14 -8x - 10 < 14

OpenStudy (anonymous):

-8x - 10 > 14

OpenStudy (anonymous):

can you help me with 2 more please?

OpenStudy (anonymous):

possibly...lets see what you got

OpenStudy (anonymous):

The table shows the solution to the equation |2x – 4| - 3 = 3: Step 1 |2x – 4| = –3 + 3 Step 2 |2x – 4| = 0 Step 3 2x – 4 = 0 Step 4 2x = 4 Step 5 x = 2 Which is the first incorrect step? Step 1 Step 2 Step 3 Solution is correct

OpenStudy (anonymous):

I believe it is step 1. Because : |2x - 4| - 3 = 3 |2x - 4| = 3 + 3 (not 3 - 3)

OpenStudy (anonymous):

oops...not -3 + 3

OpenStudy (anonymous):

The length of a rectangle is 24 units. Can the perimeter x of the rectangle be 60 units when its width y is 11 units? No, the rectangle cannot have x = 60 and y = 11 because x = 48 + 2y. No, the rectangle cannot have x = 60 and y = 11 because x = 24 + 2y. Yes, the rectangle can have x = 60 and y = 11 because x + y is greater than 48. Yes, the rectangle can have x = 60 and y = 11 because x + y is greater than 24.

OpenStudy (anonymous):

P = 2L + 2W 60 = 2(24) + 2(11) 60 = 48 + 22 60 = 70 (incorrect) first answer choice

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

no problem..glad I could help :)